solve by factoring 6x^2 -21x -12= 0
Answer by edjones(8007)

6x^2 -21x -12= 0
6*-12=-72, Two factors of -72 whose sum is -21 are -24 and +3
6x^2+3x-24x-12=0
3x(2x+1)-12(2x+1)=0
(3x-12)(2x+1)=0
3x=12, 2x=-1
x=4, x=-1/2
.
Ed